### v1

# BitMark for Infinity: Watermarking Bitwise Autoregressive Image Generative Models

Abstract
State-of-the-art text-to-image models like Infinity generate photorealistic images at an unprecedented speed. These models operate in a bitwise autoregressive manner over a discrete set of tokens that is practically infinite in size. However, their impressive generative power comes with a growing risk: as their outputs increasingly populate the Internet, they are likely to be scraped and reused as training data—potentially by the very same models. This phenomenon has been shown to lead to model collapse, where repeated training on generated content, especially from the models’ own previous versions, causes a gradual degradation in performance. A promising mitigation strategy is watermarking, which embeds human-imperceptible yet detectable signals into generated images—enabling the identification of generated content. In this work, we introduce BitMark, a robust bitwise watermarking framework for Infinity. Our method embeds a watermark directly at the bit level of the token stream across multiple scales (also referred to as resolutions) during Infinity’s image generation process. Our bitwise watermark subtly influences the bits to preserve visual fidelity and generation speed while remaining robust against a spectrum of removal techniques. Furthermore, it exhibits high radioactivity, i.e., when watermarked generated images are used to train another image generative model, this second model’s outputs will also carry the watermark. The radioactive traces remain detectable even when only fine-tuning diffusion or image autoregressive models on images watermarked with our BitMark. Overall, our approach provides a principled step toward preventing model collapse in image generative models by enabling reliable detection of generated outputs.
1 Introduction
The image generation capabilities of large scale models [7, 10, 21, 24] have improved tremendously. These models can now generate at an unprecedented speed (of less than a second) high-quality photorealistic outputs, which are often indistinguishable from real images [10]. One of the current state-of-the-art image generative models is Infinity [10], which operates in a bitwise autoregressive manner [32, 37] over an almost infinite number of tokens and yields high-resolution images. This performance, however, comes at the expense of a training process that is highly time-, cost- and most of all data-intensive [23, 27]. The required training data are often scraped from the Internet [11, 22, 30]. This practice carries risks: with the ever-growing amount of generated data on the internet, the models get increasingly trained on generated data rather than natural data. The iterative training of models on their own generated data has been shown to degrade performance [1, 31], which is referred to as model collapse. The process also leads to encoding the current models’ mistakes, biases, and unfairnesses into its future versions [36]. From the perspective of model owners, especially big companies such as OpenAI, Google, or ByteDance, who invest in large scale proprietary models that are exposed via public APIs, this poses a significant threat. To prevent the possible degradation of model quality, it is crucial for them to identify their own generated content.
A promising way to address the above problems is watermarking [4, 8, 9, 12, 20, 34, 43]. By embedding human-imperceptible, yet detectable signals into generated images, a model owner can automatically identify and filter out their model’s previously generated content when collecting training data for future versions. To prevent model collapse, watermarks should be radioactive [6, 26, 29]—that is, they should persist across model lineages, so that models trained on watermarked outputs also reproduce the watermark. This is crucial when third parties fine-tune models on such outputs and use them to generate new content, which may later be scraped and reused by the original model—causing it to (unknowingly) train on data derived from its own outputs. While watermarks have been extensively studied in diffusion models (DMs) [8, 9, 34, 43], no watermarking methods have been proposed for image autoregressive models such as Infinity, so far. Additionally, existing DM watermarks have been shown to lack radioactiveness [6], motivating the need for a new approach.
To close this gap and provide a robust way of detecting images generated by state-of-the-art models like Infinity [10], we propose the BitMark watermark that aligns with Infinity’s bitwise image generation process and operates directly on a bit-level. We show that this bit-level design is crucial in image autoregressive models, because watermarks at the token level—a common practice in autoregressive language models [14, 41]—are suboptimal. This is because unlike language models, which operate on discrete tokens and yield stable token representations, image models must decode to and re-encode from a continuous image space, introducing discrepancies between the original and recovered tokens. These discrepancies can decrease the watermark signal. Yet, we find that the discrepancies are substantially smaller at the bit level (Figure 1), motivating our bitwise approach to obtain a robust watermark. To embed the watermark, BitMark first divides all possible n-bit sequences into green and red lists, where the green list is supposed to contain sequences indicating the presence of the watermark. The embedding process of the watermark aims at completing one of the green-list bit sequences for every next bit by slightly increasing each logit of the preferred bit that would complete the sequence. This eventually causes the generated image to contain more sequences from the green list than natural images, which can be used for detection. Concretely, the detection of the watermark is performed on re-encoded images and based on the count of bit sequences from the green list.
Our watermark exhibits many desirable properties. Most importantly, the generated images are of the same quality as images generated without the embedded watermark. Moreover, the watermark embedding has a negligible impact on the speed of the image generation. Also, the corresponding watermark detection method is fast and reliable. As in natural images all bit sequences are equally likely to occur while our watermarked images have the bias towards sequences from the green list, natural images are extremely unlikely to be detected as watermarked. Furthermore, working on the bit-level makes BitMark robust against a wide range of attacks, such as adding Gaussian noise or blur to the image, applying color jitter, random crop, or JPEG compression. Our watermark is also highly resilient against dedicated watermark-removal techniques, for example, the watermark in the sand attack [38] or CtrlRegen [18]. Finally, we show that BitMark is radioactive. This means that when we use the watermarked images to train or fine-tune other image generative models, our watermark is still present in the outputs from these other models. We demonstrate all these properties through our extensive experimental evaluation.
In summary we make the following contributions:
  • We propose BitMark—the first watermarking scheme for Infinity. BitMark leverages Infinity’s unique structure and bitwise image generation process to embed the watermark on the level of bits, which preserves the high quality of the models’ outputs and its inference speed. Our approach can be applied at inference time and does not require additional retraining.
  • We thoroughly evaluate BitMark against a wide variety of standard and dedicated watermark removal attacks and show its robustness to all of them. Specifically, BitMark is also resilient to the watermark in the sand attack [38], which is targeted at removing the watermarks from the generated content. Finally, BitMark is also robust against a Bit-Flipper attack carefully crafted against it.
  • We also assess BitMark’s radioactivity and show that when other image-generative models are trained or fine-tuned on images watermarked with our method, these new models’ outputs will also contain our watermark. This makes BitMark highly suitable to identify generated content and contributes to preventing model collapse due to iterative training a model on its generated images.
2 Background
We consider the image autoregressive models, with the focus on the state-of-the-art text-to-image Infinity model that operates in a autoregressive manner, and the techniques used to robustly watermark image generative models.
Visual AutoRegressive Modeling (VAR) [32]. VAR leverages discrete quantization to predict tokens as residual maps, scaling from an initial low scale (coarse-grain) to a final scale (fine-grain) resolution. VAR further leverages the transformer architecture [33] across all tokens to significantly improve generation quality. VAR was further extended from class-to-image to the Infinity model [10], which is a text-to-image autoregressive model for high-resolution image synthesis, using next-scale prediction similar to VAR [32]. Instead of leveraging index-wise tokens for quantization like VAR [32], Infinity uses bitwise token prediction with an implicit codebook and applies binary spherical quantization [42] to select an entry from the codebook. This results in more efficient computation and further enables the usage of large codebook sizes, e.g., $2^{64}$. Random bit flipping further empowers the model to self-correct on each scale, leveling out errors stemming from earlier tokens. The visual autoregressive paradigm instantiated by Infinity achieves the state-of-the-art (SOTA) in image generation.
Watermarking Images. Watermarking entails embedding a message, pattern, or information into an image, that can be detected. The embedded information should be hard to remove and can be also used to verify the provenance of the image. In general, the watermarking methods can be either static or learning-based. In static watermark embedding, a transform is directly applied to an image [20]. Learning-based methods are generally based on three key components: a watermark (w), an encoder (E) and a decoder (D). The encoder is trained to embed the watermark into the image, while not impacting its quality and the decoder is trained to reconstruct the watermark given a watermarked image to perform detection. An important aspect of watermarks is the trade-off between their detectability and their impact on the quality of the data, which was thoroughly analyzed in [35].
Watermarking Diffusion Models. Learning-based watermarks represent the SOTA for Diffusion Models [8, 9, 34, 43]. They target different points of the image generation process, such as the initial data used to train the model [43], noise prediction [9, 34], or the decoding process [8]. The Recipe method [43] embeds a binary signature into training data using a pretrained encoder-decoder and trains an explicit decoder to extract this watermark from outputs. In contrast, Tree-Ring [34] embeds a structured pattern in the initial noise vector, designed in the Fourier space, achieving robustness to transformations by enabling detection through inversion of the diffusion process. On the other hand, the PRC Watermarking [9] generates the initial noise vector from a pseudorandom error-correcting code, allowing deterministic reconstruction of the noise for watermark retrieval. Finally, Stable Signature [8] fine-tunes the latent decoder of latent diffusion models (LDMs) to insert a recoverable watermark using a pretrained extractor, modifying the generation process at the decoding stage. However, the watermarking techniques for diffusion models are not radioactive [6], hence, do not prevent the model collapse.
Watermarking Autoregressive Models. As of now, all the efforts on watermarking autoregressive models have been directed towards creating watermarks for Large Language Models (LLMs) [14]. The KGW watermark [14] uses the concept of a green and red token lists, where generation is softly nudged to predict tokens from the green list rather than from the red list. The lists are generated dynamically, based on the hash value of the previously predicted token(s). For detection, the LLM used to generate the text is not needed. Instead, a suspect text is assessed based on how often tokens from the green list appear, compared to tokens from the red list. At the end of detection a statistical test is performed. KGW was extended to the Unigram-Watermark [41], which simplifies the design by introducing a fixed global green and red lists. The simplified approached allowed for guaranteed generation quality, provable correctness in watermark detection, and robustness against text editing and paraphrasing. In this work, we we build on these watermarking schemes originally developed for large language models (LLMs) to address the lack of watermarking techniques for image autoregressive models, with focus on the state-of-the-art Infinity model.
Private Watermarking. In our watermark, we follow the Private Watermarking from KGW [14]. For this end, the generative model is kept behind a secure API with a secret key $\mathcal{K}$. This secret key $\mathcal{K}$ is used as input to a pseudorandom function (PRF) $F$, with the $h$ previously generated tokens, to compute the red-green list split for the next predicted token. Here the context window $h$ is a hyperparameter, which trades robustness with privacy. For small context windows $h$ (e.g., 1 or 2) the watermark is robust against removal attacks but an attacker could tabulate the frequency of token pairs and brute-force knowledge about the green list, reducing the privacy of the method. For larger values of $h$ (e.g., 5), changing a single token has an impact on the next $h$ downstream tokens, potentially flipping them to the red set. The privacy of this approach can also be boosted by using $k$ different private keys and choosing one randomly during generation.
Watermark Robustness. To reliably perform image provenance, a watermark must be robust to removal attacks. There is a wide range of these kinds of attacks that have been proposed in the literature [2, 18, 25]. For example, the scrubbing attacks aim at removing the watermarks from the outputs of a generative model, which can be done by producing a valid paraphrase that is detected as non-watermarked text [13] or image [3]. Zhang et al. [38] claim that under the assumption that both a quality oracle (that can evaluate whether a candidate output is a high-quality response to a prompt) and a perturbation oracle (which can modify an output with a nontrivial probability of maintaining quality) exist, it is feasible to remove existing watermarks in generative models without quality degradation of the watermarked content. We evaluate our BitMark against a wide spectrum of attacks and show that our method is robust against them.
3 BitMark Method
The Infinity architecture [10] follows an autoregressive generation paradigm, but instead of directly predicting tokens like other autoregressive models, such as VAR [32] or RAR [37], it predicts bits. We leverage this prediction scheme to design a dedicated watermark. Our BitMark watermark is embedded across all scales (resolutions) of Infinity, which subtly shift the bitwise representation towards a higher frequency of detectable patterns to resist watermark removal and manipulation.
Problem Formulation and Desiderata. Our goal is to embed a watermark during Infinity’s image generation process and subsequently detect it by analyzing the generated image. BitMark targets the prevention of model collapse. In this setup, only the model owner is concerned with the watermark and wants to ensure that their own data are not used to train their next models. Thus, we aim at four essential properties for our watermarking scheme to prevent model collapse: (1) Negligible Impact on the Generation. To maintain the utility of the generative model, the watermark should not degrade the quality of the generated images, while also maintaining a fast inference speed. (2) Reliable Verification. The model owner should be able to detect the watermarked images and verify the confidence of the extracted watermark. The model owner encodes the image to their own space and it is only relevant to the model owner if the watermark is present there or not. The encoder part of the model is also kept private so only the model owner can verify the watermark. The natural images should not be detected as containing the watermark. At the scale of the large models trained on trillions of samples, a small number of false positives is acceptable as long as it does not exceeds a certain threshold (e.g., 1%), which can also prevent model poisoning [40]. (3) Robustness. The watermark should be robust not only against conventional attacks, such as Gaussian noise or JPEG compression, but also against strong dedicated watermark removal attacks such as CtrlRegen [18] or watermarks in the sand [38]. (4) Radioactivity. If other generative models are trained on watermarked images, images generated by these new models should also contain our watermarks [6, 26, 29].
Threat Model. We assume two parties: a model owner providing an image generation API where our BitMark is deployed, and an attacker who aims to fool the model owner to train on their own generated images and to prevent the use of large-scale natural image datasets for model training. We assume that the model owner follows the Private Watermarking from [14] and keeps the generative model behind a secure API. The model owner wants to watermark their own images in order not to train on them. The watermark should also be radioactive so that if any other party fine-tunes their own models on the watermarked data, the fine-tuned model’s outputs are watermarked, as well. Further, we assume that the watermarking technique is private and verifiable by the model owner. On the other hand, the attacker might attempt to remove the embedded watermark. For instance, competitors might still want to remove the watermark to harm the model owner. We assume that the attacker has black-box access to the model, i.e., they can provide inputs and observe the corresponding outputs. The attacker is allowed to make arbitrary modifications to the generated image, including but not limited to: applying image transformations (e.g., noising, blurring, cropping, etc.) or specific watermark removal techniques [18, 38]. We define adversarial behavior as any attempt to remove the watermark in the generated image while preserving high perceptual quality and semantic content. Attacks that destroy the usefulness or visual coherence of the image (e.g., replacing it with noise) are not considered meaningful threats.
3.1 Motivation for the Bitwise Watermark
We identify an inherent challenge in watermarking the images generated by autoregressive models like Infinity, stemming from imperfections in image encoding and decoding. Specifically, the model first generates a sequence of tokens, which are decoded into an image. However, when the exact same image is re-encoded and tokenized, it produces a different set of tokens.
Figure 1: Bit and token overlaps in Infinity.
This also holds partly true for bits that underlie the tokens. The observed behavior does not occur in the autoregressive LLMs, where the tokens produced by decoding and then re-encoding a generated text match exactly. We quantify the bit and token discrepancies in Figure 1. In both cases, bits and tokens are generated by Infinity and then reconstructed by its auto-encoder. For Infinity, the highest token overlap occurs at the initial and final resolutions (scales). This is likely because the early stages capture coarse structural features, while the final resolution refines the image, leading to fewer changes in the token sequence. We observe a consistent average token overlap of approximately 2.385% and the bit overlap of 77.432%. Given the significantly large codebook size of $2^{n}$ (where $n\in\{32,48,64\}$) in Infinity, we observe that the bit-wise information loss is much smaller than the token-wise loss. This motivates our work to design a watermark that biases the generated sequences of tokens on the bit-level instead of on the token-level.
3.2 Embedding our BitMark Watermark
Next, we detail how BitMark is embedded. The embedding protocol is directly applied on the bit-level in the autoregressive generation process of the Infinity model. We present the full procedure in Algorithm 1. First, we separate the list of all bit sequences of length $n$: $S=\{0,1\}^{n}$ into a red list $R$ and a green list $G$. Intuitively, we softly nudge the model to generate bit sequences of length $n$ that are part of the green list $G$. For each next bit prediction $b_{j}$, if the bit can complete one of the bit sequences from $G$, we add a small bias $\delta$ to the logit $l_{j}^{(b_{j})}$ of the bit $b_{j}$:
$p_{j}=\begin{cases}\frac{\exp(l^{(b_{j})}_{j}+\delta)}{\exp(l^{(\neg b_{j})}_{j})+\exp(l^{(b_{j})}_{j}+\delta)},\quad pre+b_{j}\in G\\ \frac{\exp(l^{(\neg b_{j})}_{j})}{\exp(l^{(\neg b_{j})}_{j})+\exp(l^{(b_{j})}_{j}+\delta)},\quad pre+\neg b_{j}\in R,\end{cases}$ (1)
where $pre=(b_{j-(n-1)},\dots,b_{j-1})$ is the prefix bit sequence of length $(n-1)$, i.e., the sequence of bits preceding $b_{j}$, $+$ is the bit concatenation operation, and $l_{j}$ is the vector of logits for the bit $b_{j}$. We apply this procedure to all bits and across all resolutions (scales) that are generated by the model. Thus, as many bit sequences as possible are biased towards $G$ and the final green fraction, i.e., the number of sequences part of the green set, of the generated bits is significantly above the threshold of 50%, the standard fraction for clean images (see Appendix D). By biasing the logit for the preferred bit, BitMark has negligible impact on image quality, as mostly high entropy bits are flipped (see Table 1). Since high entropy bits have sufficiently small difference in the probability of assigning them to value 00 or $1$, the relatively small added bias influences the final value. Additionally, biasing only the logits makes the watermarking process difficult to revert, as it is difficult to know for an attacker which bits are of high entropy and can be flipped to the red list $R$ without degrading the final image quality.
3.3 Watermark Detection
Watermark detection is performed directly on the bits of encoded images. All that is required is knowledge about the green and red lists, as well as access to the tokenizer. Intuitively, we count how often each sequence of bits in the encoded image follows the green list and compare how the frequency differs from non-watermarked images.
Detecting the Watermark. The detection protocol follows the standard encoding from Infinity and is detailed in Algorithm 2. First we obtain the bit representations for each resolution. We then use our knowledge of $G$ and $R$ to test how often the bit sequences from the red and green lists appear and count them. This results in the knowledge of how many green sequences appear in the encoded image, which is random for every non-watermarked image, i.e., $50\%$ (see Appendix D).
Robust Statistical Testing. Given the final number of green and red sequences detected in the encoded image, we perform a statistical test based on the null hypothesis $\mathcal{H}_{0}:$ The sequence of bits is generated with no knowledge about the green and red lists. Intuitively, this assumes that a given image is not watermarked. A watermarked image consists of $\gamma T$ green and $(1-\gamma)T$ red bit sequences, where $\gamma$ determines the size of the green list and $T$ is the total number of bits in the encoded image that our watermark can influence. Per a given token $t$ with $m$ bits, BitMark influences $k=m-(n-1)$ bits, since we start watermarking from the first $n$ bit pattern. Thus, overall we have $T=\sum_{i=1}^{K}r_{i}\cdot k$, where $r_{i}$ is number of tokens per resolution $i$. A natural image is expected to violate the red list rule with the half of its bits, while a watermarked image violates this rule less frequently. If the null hypothesis is true, then the number of green list bit sequences denoted C has the expected value of $\gamma T$ and variance $\gamma(1-\gamma)T$. To ensure statistical validity, we follow [14] and use a z-test. We reject the null hypothesis and detect the watermark if $z$ is above a chosen threshold:
$z=(C-\gamma T)/\sqrt{T\gamma(1-\gamma)}.$ (2)
Algorithm 1  Watermark Embedding
Inputs: prompt $x$, tokenizer $tok$, autoregressive model $f_{\theta}$ with parameters $\theta$, green set $G$, red list $R$, image decoder $\mathcal{D}$, BSQ quantizer $\mathcal{Q}$.
Hyperparameters: steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of tokens for resolution $i$ is $r_{i}$, constant $\delta$ added to the logits of the $green$ list bits, $n$ the length of the bit sequences in the lists ($G$ and $R$), the number of bits per token $d$.
$e=tok(x)$
for $i=1,\dots,K$ do
       $(l_{1},\dots,l_{r_{i}\cdot d})=f_{\theta}(e,h_{i},w_{i})$for $t=1,\dots,r_{i}$ do
             for $j=n,\dots,d$ do
[…]
      $u_{i}=(b_{1},\dots,b_{r_{i}\cdot d})$$z_{i}=\mathcal{Q}(u_{i})$ $z_{i}=\text{upsample}(z_{i},h_{K},w_{K})$$e=e+z_{i}$
$im=\mathcal{D}(e)$
Return: watermarked image $im$
Algorithm 2  Watermark Detection
Inputs: raw image $im$, green list $G$, red list $R$, image encoder $\mathcal{E}$, BSQ quantizer $\mathcal{Q}$
Hyperparameters: steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of tokens for resolution $i$ is $r_{i}$, the number of bits per token $d$.
$e=\mathcal{E}(im)$
$C=0$
for $i=1,\dots,K$ do
       $u_{i}=\mathcal{Q}(\text{downsample}(e,h_{i},w_{i}))$ $u_{i}=(b_{1},\dots,b_{r_{i}\cdot d})$ $C$ = Count($(b_{1},\dots,b_{r_{i}\cdot d})$, $G$) $u_{i}=\text{upsample}(u_{i},h_{K},w_{K})$$e=e-u_{i}$
Return: StatisticalTest($\mathcal{H}_{0},(C)$)
4 Empirical Evaluation
Experimental Setup. We employ the Infinity-2B checkpoints and code  [10] to evaluate BitMark across different watermarking strengths and against a wide variety of attacks.  [^1] Infinity-2B was released six months ago, at the time of submission, allowing sufficient time for experimentation. Note that the most powerful Infinity-20B model (that we are aware of) is kept private (while the Infinity-8B was released less than 3 months ago before the submission, limiting our ability to test it). The Infinity-2B model generates $1024\times 1024$ pixel RGB-images and we choose the visual-tokenizer with a vocabulary size of $2^{32}$ per token. We follow [10] in the choice of hyper-parameters and use a classifier-free-guidance of 4, with a generation over 13 scales, following the scale schedule for an aspect ratio of 1:1. Hence, we utilize the best performing model, tokenizer and hyperparameters reported in [10]. We ablate the choice of the green, red list split for our experiments in Appendix B.
4.1 BitMark Introduces Detectable Watermarks while Maintaining Generation Performance
We demonstrate that our BitMark preserves the utility of the Infinity model with negligible impact on the image quality and the inference speed.
High Image Quality. Table 1 presents a quantitative evaluation of image generation quality for different watermarking strengths dependent on the logit bias parameter $\delta$. We generated 10,000 images based on prompts of the MS-COCO 2014 [16] validation dataset. We report the FID, KID, measuring the similarity of generated images and natural images in feature space, and the CLIP Score, which measures the alignment between prompt and image, for the varying watermark strengths $\delta$.
Table 1: Image quality mean (std) vs $\delta$.
[…]
0 33.36 1.42 (0.12) 31.16 (0.28)
1 32.61 1.38 (0.13) 31.16 (0.28)
2 31.05 1.26 (0.12) 31.15 (0.28)
3 29.61 1.03 (0.08) 31.03 (0.27)
4 42.98 1.78 (0.08) 29.78 (0.32)
5 127.44 11.16 (0.34) 26.13 (0.40)
The results show that, for $\delta\leq 3$, BitMark has a negligible impact on image quality (as measured by FID and KID) and minimally affects the prompt-following capabilities of the underlying model, as indicated by the CLIP Score. We also present a qualitative analysis of image quality in Section E.2, which shows images generated for different $\delta$. While there is an inherent trade-off in terms of watermark strength an image quality, we choose a small $\delta=2$ for our watermarks to obtain minimal impact on the generation while having a significantly strong watermark.
Fast Watermark Generation and Verification. For a watermark to be reasonably applicable in a real world setting, the watermark should have negligible impact on image generation speed [8, 34, 43]. We thus analyze the practicality of applying BitMark in a standard setting and its impact on generation speed of the Infinity model. Next to the generation performance, the speed of watermark detection has a strong impact on the possible applications, especially in a setting where we want to quickly detect data generated by Infinity. The quantitative analysis of BitMark in Section E.3 shows that generating a watermarked image requires at most one-tenth the time longer than the standard image generation. Verification, i.e., encoding an image into its bit representation and computing the $z$-score, takes less than 0.5 seconds per image (usually around 5X faster than the image generation).
4.2 BitMark is Robust Against a Wide Range of Attacks
Our results demonstrate that BitMark is extremely robust against a wide range of attacks, including conventional attacks, reconstruction attacks, watermarks in the sand attack, and even our own adaptive bit-flipping attack, which we carefully design with the aim of exploiting knowledge of our watermark method to remove BitMark.
Table 2: Rotation attack.
Degrees TPR@1%FPR $z$ (std)
-20, 20 99.9 14.87 (9.77)
-30, 30 87.4 11.78 (9.55)
-40, 40 70.0 9.66 (9.38)
-50, 50 56.8 8.21 (8.99)
Robustness against Conventional Attacks. We perform a wide range of conventional attacks, namely: 1) Noise: Adds Gaussian noise with a std of 0.1 to the image, 2) Blur Noise attack + application of a Gaussian blur with a kernel size of 7, 3) Color: Color jitter with a random hue of 0.3, saturation scaling of 3.0 and contrast of 3.0, 4) Crop: Random crop to 70% 5) Rotation: Random rotation between different degrees in Table 2 and 0° and 180° in Table 3. 5) JPEG: 25% JPEG compression. Table 3 shows that our approach achieves a high TPR@1%FPR for most conventional attacks. Furthermore, BitMark achieves a high TPR@1%FPR against most attacks even for a low watermarking strength $\delta=1$. Although BitMark has slightly lower performance towards arbitrary rotation angles, the watermark is still highly detectable and it still obtains high TPR@1%FPR within a given rotation threshold. We report TPR@1%FPR and mean (std) of $z$ for different degrees of rotations for 1,000 images in Section E.5. Notably, it achieves as high as 99.9% TPR@1%FPR when the rotation range is between -20° and 20° Table 2. Given the strength of off-the-shelf tools for rotation estimation [19], it is highly achievable to rotate the image back into a tolerable range of rotations during the pre-processing stage in the model training process, thus ensuring that the watermark is robust against image rotations.
Robustness against Reconstruction Attacks. We also evaluate our proposed method against two state-of-the-art reconstruction attacks (shown as the last two columns in Table 3): 1) SD2.1-VAE: Encoding and decoding with the Stable Diffusion 2.1 autoencoder [24], 2) CtrlRegen+: Application of controllable regeneration from noise in the latent space following  [18]. For the reconstruction attacks, we find that the strength $\delta=2$ can withstand the CtrlRegen+ reconstruction attack, which removes watermarks such as TreeRing [34] or StableSignature [8].
Table 3: BitMark is robust against watermark removal attacks. We report the TPR@1%FPR (%) for the different attacks.
$\delta$ Conventional Attacks Reconstruction Attacks
None Noise Blur Color Crop Rotation JPEG SD2.1-VAE CtrlRegen+
1 100.0 97.8 99.2 99.3 68.5 14.5 100.0 100.0 61.9
2 100.0 99.6 99.9 99.8 98.8 20.1 100.0 100.0 91.6
3 100.0 99.7 99.9 99.9 99.8 23.8 100.0 100.0 97.0
4 100.0 99.9 99.9 100.0 99.9 29.2 100.0 100.0 99.0
5 100.0 100.0 100.0 99.9 100.0 34.3 100.0 100.0 99.8
Table 4: BitMark is robust against the watermarks in the sand attack [38]. We present the detection results of our watermarking scheme ($\delta=2$) before and after the attack. We use 200 iterations, and generate 100 candidates during each iteration.
Setting Accuracy AUC TPR@1%FPR Green Fraction $z$ $z$ Threshold
(%) (%) (%) (%) @1%FPR
Original Image - - - 49.7 0.2 -
Watermarked Image, before the attack 100.0 100.0 100.0 57.6 87.8 59.6
Watermarked Image, after the attack 98.5 99.8 89.0 50.9 10.8 6.8
Robustness against Watermarks in the Sand Attack. We also evaluate the robustness of our watermarking scheme against the attack proposed in [38], namely, the watermarks in the sand attack. The attack leverages the CLIP Score as a quality oracle, the stable-diffusion-2-base [21] as a perturbation oracle, and iteratively modifies the watermarked image to remove the watermark. During each iteration, the attack produces potential candidates by inpainting randomly masked regions in the image. The attack successfully replaces most visual details in the original watermarked image while selecting the candidate with the best CLIP Score. As a result, many watermarks for diffusion models (e.g., [8]) can be removed through this attack. However, results in Table 4 show that our proposed method achieves a significant detection rate with TPR@1%FPR of 89.0%. Although a large percentage of the green-list patterns are removed, the average $z$-score in the attacked images is still as high as 10.8, which remains significant when compared with a mean $z$-score of 0.2 for non-watermarked images. Moreover, the attack causes a quality loss of the image when compared to the original image generated by Infinity. With 200 iterations and 100 candidates during each iteration, the method further requires a high computational cost, where more than 12 minutes are needed to attack a single image in the environment specified in Section C.2. We note that, the results do not conclude the failure of the theoretical framework in watermarks in the sand, but rather that there are currently no successful instantiations of the perturbation oracles and quality oracle for Infinity that can remove our proposed watermark. A more detailed analysis about this attack can be found in Appendix F.
Figure 2: Bit-Flipper attack for $\delta=2$
Robustness Against our Novel Bit-Flipper Attack. We further demonstrate that, even giving the attacker much stronger knowledge than our threat model formulated in Section 3, BitMark cannot be erased with severe image degradation. Concretely, we assume the attacker is given the knowledge of the auto-encoder of the generative model, the watermarking embedding and detection algorithms, and even our red and green lists. We propose a novel Bit-Flipper attack tailored for this broader attack surface, aiming at removing our watermarks. During this attack, the attacker first computes the total number of bit sequences $|s|$ and calculate the green fraction $C/|s|$. Then, the attacker randomly flips bits to convert green bit sequences to red ones. The goal of the bit flipping is to reduce the green fraction to the level of non-watermarked images, which is 50%, as shown in Appendix D. Specifically, the attacker computes the probability of flipping a bit via a flip factor $\phi$, such that $p=($$C/|s|$$)-0.5)*\phi$, whereas $\phi$ determines the bit sequence removal severity. A detailed algorithm of our attack is described in Algorithm 3. As shown in Appendix G, $\phi=2.2$ has the most impact on detectability. However, randomly flipping bits comes with very high degradation in image quality, as shown in Section E.2. Considering a symmetric detection threshold $|z|$, if we remove too many bit sequences in $G$, this further leads to the detectability of the watermark (see Appendix G).
4.3 Our BitMark Watermark Exhibits High Radioactivity
We demonstrate that our BitMark exhibits high radioactivity, i.e., other generative models fine-tuned on our watermarked images also produce our bit patterns that are reliably detectable.
Formalizing Radioactivity. We denote by model $M_{1}$ our original model that generates watermarked outputs with a watermarking method, e.g., our BitMark. These watermarked outputs might be used as a training set for a second model $M_{2}$. In the context of language models it has been shown, that there are watermarks that persist this process and remain detectable in the outputs of $M_{2}$, a property referred to as radioactivity [29]. This property enables the detection of generated content, even when it has been used to train subsequent models. However, recent findings indicate that this property does not hold for watermarks in image generative models, such as diffusion models [6].
Training Setup. Given the extensive compute costs of training high-performant image-autoregressive models from scratch, we simulate full training by full fine-tuning the models. Therefore, we first generate 1,000 images with prompts of the MS-COCO2014 [16] validation dataset with our BitMark embedded under a watermarking strength of $\delta=2$. The watermarked outputs of $M_{1}$ are then used to train (i.e., in our approximative setup full fine-tune) the $M_{2}$ model. We analyze two different settings, with the same $M_{1}$ of Infinity-2B, namely: (1) the same modality setting, transfer to Infinity-2B and (2) the cross-modality setting, transfer to the Latent Diffusion Model (LDM) Stable Diffusion 2.1 [24]. To test the radioactivity, we compute the TPR@1%FPR of output images of $M_{2}$ and report them in Table 5.
Table 5: BitMark exhibits strong radioactivity. We report the TPR@1%FPR (%).
Type of $M_{1}$ Type of $M_{2}$ Output of $M_{1}$ Output of $M_{2}$
Infinity-2B Infinity-2B 100.0 99.7
Infinity-2B Stable Diffusion 2.1 100.0 98.9
BitMark is Highly Radioactive. Table 5 highlights the strong radioactivity of our BitMark. When fine-tuning a model $M_{2}$ of the same architecture, the watermark is still detectable with a 99.7% TPR@1%FPR. Even in the cross-modality setting, BitMark is still detectable with a 98.9%TPR@1%FPR. Dubiński et al. [6] attribute the lack of radioactivity of watermarks in LDMs to the encoding from the pixel space into the latent space and the noising-denoising process while training $M_{2}$. However, our BitMark remains detectable despite the encoding and the noising-denoising process. The robustness of our watermark arises from slightly changing all parts of a given generated image, including low-level and high-level features, which are then also partly adapted by $M_{2}$. The observed results are in line with our robustness against CtrlRegen+ (see Table 3), an attack which simulates noising and denoising in the latent space. BitMark’s radioactivity is a very useful property that allows model owners to track provenance of their data and data generated based on it, even when this new data does not originate directly from their original model.
5 Conclusions
We present BitMark, the first bitwise watermarking framework for image autoregressive models, such as Infinity. We show that our watermark can be efficiently embedded and detected, imposes negligible overhead on image quality and generation speed, while exhibiting a strong robustness against a wide range of watermark removal attacks, including the strong watermarks in the sand and our own adaptive Bit-Flipper attack. Moreover, BitMark exhibits strong radioactivity, i.e., downstream models fine-tuned on our watermarked outputs consistently inherit and reproduce our bit patterns, which makes our watermark reliably detectable. This is important for model owners to identify images derived from their own generated images, and exclude them from training of future model versions. Overall, BitMark not only sets a new standard for watermarking in autoregressive image generation, but also provides a critical defense against model collapse, contributing to further advancements in image generative models.
Acknowledgments and Disclosure of Funding
This work was supported by the German Research Foundation (DFG) within the framework of the Weave Programme under the project titled "Protecting Creativity: On the Way to Safe Generative Models" with number 545047250. Responsibility for the content of this publication lies with the authors.
Appendix A Additional Algorithms
We report the functions used in Algorithm 1, Algorithm 2 and Algorithm 3 in detail here:
  • Bias: Addition of the bias $\delta$ to the current logit $l_{c}$ based on the green list $G$.
  • Sample: Use the probabilities in $p_{c}$ to sample the returned bit, either 0 or 1.
  • lookup: Perform the BSQ lookup [42] to return the quantized states.
  • interpolate: Bilinear scaling to target resolution.
  • Count: Action performed for the statistical test, which counts the number of bits which are part of the green list.
  • $\phi_{i}$ extra convolution operations to address the information loss (following [32]).
We present the pseudo code of the novel Bit-Flipper attack in Algorithm 3. To apply the attack, we have to first encode the image in to bits (line 1). During deconstructing the old watermarked image, we construct a new attacked image. We recursively iterate through all scales $K$, retrieve the bit representations of the respective scale (line 4) and count how many bits are part of the green list and the red list (line 5). Then we define the probability of flipping a bit based on the green fraction and the flip factor $\phi$. Next, we change the bits which are part of the green list based on the random probability $p$ (line 8, 9). We follow the standard encoding and decoding process, to iterate through all resolutions of the image (line 10-13). Finally, we return the decoded version of the attacked image.
Algorithm 3  Bit-Flipper Attack
Inputs: image $im$, green list $G$, red list $R$, image encoder $\mathcal{E}$, quantizer $\mathcal{Q}$, image decoder $\mathcal{D}$.
Hyperparameters: Flip factor $\phi$, steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of bits for resolution $i$ is $t_{i}$, $n$ the length of the bit vector, $|s|$ the total number of bit sequences.
$e=\mathcal{E}(im)$
$img=\text{init()}$()
for $i=1,\dots,K$ do
       $u_{i}=\mathcal{Q}(\text{interpolate}(e,h_{i},w_{i}))$ $C_{i}$ = Count($(b_{1},\dots,b_{t_{i}})$, $G$) $p=(\frac{C_{i}}{|s_{i}|}-0.5)*\phi$ for $j=n,\dots,t_{i}$ do
             if $(b_{j-n},\dots,b_{j})\in G\land\text{random}()<p$ then
                   $b_{j}=\lnot b_{j}$Flip bit with probability $p$ if sequence is in green list}
      $z_{i}=\text{lookup}(u_{i})$ $z_{i}=\text{interpolate}(z_{i},h_{K},w_{K})$ $e=e-\phi_{i}(z_{i})$ $img=img+\phi_{i}(z_{i})$
Return: attacked image $\mathcal{D}(img)$
Appendix B Green and Red List Selection
Selection of Red and Green Lists. The re-encoding loss presented in Section 3.1 also motivates the choice of a small length of the green list $n$, as larger consecutive bit sequences are more likely to be interrupted. Additionally, this auto-encoder has the property of re-encoding images towards an equal frequency of 0’s and 1’s, so it is not efficiently possible to bias each bit towards either 0 or 1, making the choice of $n=1$ unpractical (see Table 6). Thus we leverage the second smallest possible choice, $n=2$, resulting in 6 possible choices to separate $G$ and $R$ by keeping $|G|=|R|$.
As we have the constraint to not bias towards 0 or 1, as this type of pattern is more likely removed by the auto-encoder, and given the constraint shown that the same prefix must not be in the same list twice as discussed in Section 3.2, the most effective $G$ for BitMark are $G=\{01,10\}$ and $G=\{00,11\}$. We choose $G=\{01,10\}$ throughout all experiments.
Changing the Length of the Bit Sequence. We additionally ablate the size of the bit sequence in Table 6. For each sequence size $n$, if both lists should have the same size, we choose $\frac{2^{n}}{2}$ out of $2^{n}$ possibilities without replacement, so there are $\binom{2^{n}}{2^{n-1}}$ possibilities of constructing the $G$ and $R$. For the 2-bit sequences, flipping the green and red list has only limited impact on the final z-score, thus for the 3-bit sequences we only analyze all unique choices, i.e., we do not interchange green and red list, leaving us with 35 distinct possible green and red lists. For $n=3$, two $G$ have a similar $z$ compared to the best 2-bit sequences. These choices have the same properties as discussed above, namely (1) they equally often bias towards 0 and 1 and (2) none of the sequences share the same prefix, i.e., the biasing is always effective.
Table 6: Green and Red list selection. We analyze all possible green and red lists for $n=1$, $n=2$ and $n=3$ with $\delta=2$ and 1,000 images. We also report the fraction of 1’s that are in the bit representation during generation of the image (1s Gen) and after re-encoding the image using the image encoder (1s Enc).
Green List $G$ Red List $R$ Average Z-Score 1s Gen 1s Enc
$\{0\}$$\{1\}$ 6.18 31.50% 49.47%
$\{1\}$$\{0\}$ 2.39 64.52% 50.21%
$\{11,00\}$$\{01,10\}$ 91.35 49.74% 49.98%
$\{01,10\}$$\{00,11\}$ 90.33 49.97% 49.99%
$\{00,10\}$$\{01,11\}$ 6.52 31.85% 49.44%
$\{11,01\}$$\{00,10\}$ 2.56 63.70% 50.22%
$\{00,01\}$$\{10,11\}$ 0.25 49.92% 49.98%
$\{11,10\}$$\{00,01\}$0.25 49.92% 49.98%
$\{000,011,100,111\}$$\{001,010,101,110\}$ 88.00 49.75% 49.98%
$\{000,011,110,111\}$$\{001,010,100,101\}$ 87.91 49.75% 49.98%
$\{000,011,101,111\}$$\{001,010,100,110\}$ 67.90 49.75% 49.98%
$\{000,100,110,111\}$$\{001,010,011,101\}$ 44.95 40.92% 49.52%
$\{000,001,011,111\}$$\{010,100,101,110\}$ 32.91 58.19% 50.24%
$\{000,001,010,101\}$$\{011,100,110,111\}$ 32.85 44.09% 49.68%
$\{000,010,011,111\}$$\{001,100,101,110\}$ 29.04 40.92% 49.52%
$\{000,100,101,111\}$$\{001,010,011,110\}$ 29.04 40.92% 49.52%
$\{000,101,110,111\}$$\{001,010,011,100\}$ 26.43 40.92% 49.52%
$\{000,011,100,110\}$$\{001,010,101,111\}$ 20.11 49.75% 49.98%
$\{000,001,011,110\}$$\{010,100,101,111\}$ 10.03 58.19% 50.24%
$\{000,001,011,100\}$$\{010,101,110,111\}$ 8.72 58.19% 50.24%
$\{000,001,010,100\}$$\{011,101,110,111\}$ 6.48 44.09% 49.68%
$\{000,010,100,101\}$$\{001,011,110,111\}$ 6.20 33.07% 49.48%
$\{000,010,100,110\}$$\{001,011,101,111\}$ 5.99 33.07% 49.48%
$\{000,010,011,100\}$$\{001,101,110,111\}$ 5.55 40.92% 49.52%
$\{000,001,010,110\}$$\{011,100,101,111\}$ 3.74 44.09% 49.68%
$\{000,010,101,110\}$$\{001,011,100,111\}$ 3.21 33.07% 49.48%
$\{000,010,011,110\}$$\{001,100,101,111\}$ 2.93 40.92% 49.52%
$\{000,100,101,110\}$$\{001,010,011,111\}$ 2.93 40.92% 49.52%
$\{000,010,100,111\}$$\{001,011,101,110\}$ 2.14 33.07% 49.48%
$\{000,001,100,101\}$$\{010,011,110,111\}$ 0.25 49.92% 49.98%
$\{000,001,010,011\}$$\{100,101,110,111\}$ 0.25 49.92% 49.98%
$\{000,011,100,101\}$$\{001,010,110,111\}$ 0.11 49.75% 49.98%
$\{000,001,101,110\}$$\{010,011,100,111\}$ 0.07 49.92% 49.98%
$\{000,011,101,110\}$$\{001,010,100,111\}$ 0.02 49.75% 49.98%
$\{000,001,100,110\}$$\{010,011,101,111\}$0.05 49.92% 49.98%
$\{000,001,101,111\}$$\{010,011,100,110\}$0.26 49.92% 49.98%
$\{000,001,100,111\}$$\{010,011,101,110\}$0.37 49.92% 49.98%
$\{000,001,110,111\}$$\{010,011,100,101\}$0.55 49.92% 49.98%
$\{000,010,101,111\}$$\{001,011,100,110\}$0.65 33.07% 49.48%
$\{000,010,110,111\}$$\{001,011,100,101\}$0.85 33.07% 49.48%
$\{000,001,011,101\}$$\{010,100,110,111\}$1.45 58.19% 50.24%
$\{000,010,011,101\}$$\{001,100,110,111\}$12.98 40.92% 49.52%
$\{000,001,010,111\}$$\{011,100,101,110\}$13.39 44.09% 49.68%
Figure 3: Longer sequences are less consistent after re-encoding. We analyze the consistency of sequences of length $n$ after re-encoding over the whole image, following the setup of Figure 1.
Consistency in Longer Sequences. We analyze the overlap of bit sequences with growing size in Figure 3 which shows that with increased length of the sequence, the consistency of the pattern decreases. This has a direct impact on the choice of $n$, as it shows that larger $n$ lead to less overlap, decreasing the impact of BitMark which is already noticeable for $n=3$ (see Table 6).
Appendix C Experimental Setup
C.1 Hyperparameters
For the experiments on radioactivity, we chose 1,000 prompts from the MS-COCO 2014 validation dataset, generated the respective number of images with the $M_{1}$ Infinity model and then fine-tuned the second model $M_{2}$ for 5 epochs with a learning rate of $10^{-4}$ and a batch-size of 4 (batch-size of 1 for Infinity-2B).
Since VAR [32] and RAR [37] are class-conditional models for the ImageNet1k dataset, we first generated ImageNet fakes using Infinity-2B with the ImagenNet class name and "A photo of {class name}" as prompt. We used the same hyperparameters such as batch-size, number of epochs and learning rate consistent with the experiments on Infinity-2B and Stable Diffusion 2.1.
C.2 Environment
Our experiments are performed on Ubuntu 22.04, with Intel(R) Xeon(R) Gold 6330 CPU and NVIDIA A40 Graphics Card with 40 GB of memory.
Appendix D Natural Images
Table 7: Detection on non-watermarked samples. Mean (std) of $z$ and green fractions for 1,000 non-watermarked samples.
Dataset $z$ Green Fraction
Imagenet1k 0.6311 (1.966) 0.5005 (0.0017)
MS-COCO2014 0.6773 (1.735) 0.5006 (0.0015)
MS-COCO2014 (Infinity-2B generated) 0.3620 (1.641) 0.5003 (0.0014)
To ensure the validity of our hypothesis we compute the average number of members of the green list and $z$ for natural images from different datasets, as well as images generated for the non-watermarked Infinity model. Table 7 shows that indeed, the green fraction for non-watermarked images is 50%.
Appendix E Additional Experiments
E.1 Detailed Analysis of BitMark
Next, we analyze the low-level behavior of BitMark during image generation. Table 9 depicts information about the generation of images with Infinity, averaged over 10,000 images. It shows that the entropy is low in earlier scales, meaning the bits are chosen with high certainty as these depend more the input prompt, but large on later scales, which means sampling either of the possible values for the bit can lead to different, but both high-quality content. This insight aligns with the re-encoding loss, which keeps more bits consistent on earlier scales.
Table 10, depicts the low-level information of applying BitMark with $\delta=2$. Applying BitMark changes up to 37% of bits per scale, yet the re-encoding loss only increases by 5% compared to the non watermarked image. Table 8 further shows the correlation between different analyzed factors. As expected, there is a high correlation between changed bits and entropy, as due to the soft-biasing we mostly change bits which have low entropy.
Table 8: Correlations of different factors when applying BitMark on all scales within the Infinity generation process. Entropy, Green Fraction, Changed Bits & Re-Encoding Loss are relative values. The Re-Encoding Loss displays the number of bits changed during re-encoding, whereas Changed Bits display the number of bits changed during the generation process. 10,000 images, $\delta=2$.
Entropy Changed Bits $z$ Green Fraction
Changed Bits 0.999 - - -
$z$ 0.706 0.689 - -
Green Fraction 0.918 0.917 0.642 -
Re-Encoding Loss 0.926 0.931 0.475 0.751
Table 9: Statistics regarding the infinity generation process for 10,000 images. Entropy, Green Fraction& Re-Encoding Loss are relative values. The Re-Encoding Loss displays the number of bits changed during re-encoding. We report the mean (std) for all values.
Scale Entropy $z$ Green Fraction Re-Encoding Loss Number of
Tokens (Bits)
1 0.051 (0.037) 0.01 (0.876) 0.501 (0.079) 0.068 (0.063) 1 (32)
2 0.108 (0.044) -0.195 (1.114) 0.491 (0.049) 0.132 (0.063) 4 (128)
3 0.133 (0.034) -0.463 (1.226) 0.49 (0.027) 0.145 (0.041) 16 (512)
4 0.178 (0.029) -0.522 (1.217) 0.492 (0.018) 0.2 (0.031) 36 (1,152)
5 0.202 (0.024) -0.558 (1.199) 0.494 (0.013) 0.226 (0.024) 64 (2,048)
6 0.215 (0.021) -0.663 (1.229) 0.495 (0.009) 0.247 (0.02) 144 (4,608)
7 0.223 (0.019) -0.5 (1.266) 0.497 (0.007) 0.239 (0.017) 256 (8,192)
8 0.223 (0.015) -0.349 (1.187) 0.498 (0.005) 0.277 (0.016) 400 (12,800)
9 0.226 (0.014) -0.122 (1.173) 0.5 (0.004) 0.279 (0.014) 576 (18,432)
10 0.243 (0.014) 0.157 (1.165) 0.5 (0.003) 0.264 (0.012) 1,024 (32,768)
11 0.237 (0.012) 0.227 (1.161) 0.501 (0.003) 0.319 (0.011) 1,600 (51,200)
12 0.233 (0.012) 0.329 (1.198) 0.501 (0.002) 0.328 (0.01) 2,304 (73,728)
13 0.234 (0.014) 0.462 (1.173) 0.501 (0.002) 0.212 (0.012) 4,096 (131,072)
1-13 0.193 (0.012) 0.232 (1.644) 0.497 (0.008) 0.266 (0.009) 10,521 (336,372)
Table 10: Statistics regarding the infinity generation process with BitMark applied on all scales for 10,000 images, $\delta=2$. Entropy, Green Fraction, Re-Encoding Loss & Changed Bits are relative values. The Re-Encoding Loss displays the number of bits changed during re-encoding, whereas Changed Bits display the number of bits changed during the generation process. We report the mean (std) for all values.
Scale Entropy $z$ Green Fraction Re-Encoding Loss Changed Bits
1 0.051 (0.037) 0.218 (0.897) 0.52 (0.081) 0.088 (0.076) 0.09 (0.077)
2 0.108 (0.043) 0.854 (1.105) 0.538 (0.049) 0.181 (0.079) 0.179 (0.075)
3 0.132 (0.032) 2.37 (1.452) 0.552 (0.032) 0.2 (0.053) 0.208 (0.05)
4 0.175 (0.028) 3.965 (1.613) 0.558 (0.024) 0.26 (0.037) 0.279 (0.042)
5 0.2 (0.024) 6.252 (1.812) 0.569 (0.02) 0.286 (0.029) 0.313 (0.035)
6 0.213 (0.021) 10.118 (2.304) 0.575 (0.017) 0.305 (0.022) 0.329 (0.029)
7 0.224 (0.019) 15.701 (2.832) 0.587 (0.016) 0.3 (0.02) 0.346 (0.027)
8 0.224 (0.016) 16.247 (2.788) 0.572 (0.012) 0.329 (0.016) 0.345 (0.022)
9 0.228 (0.015) 20.876 (3.341) 0.577 (0.012) 0.333 (0.013) 0.351 (0.02)
10 0.244 (0.015) 32.837 (5.104) 0.591 (0.014) 0.319 (0.011) 0.373 (0.019)
11 0.237 (0.013) 29.535 (5.06) 0.565 (0.011) 0.361 (0.009) 0.363 (0.017)
12 0.242 (0.013) 34.973 (6.681) 0.564 (0.012) 0.37 (0.008) 0.368 (0.017)
13 0.247 (0.014) 64.333 (11.823) 0.589 (0.016) 0.256 (0.015) 0.374 (0.019)
1-13 0.194 (0.012) 90.784 (14.852) 0.566 (0.013) 0.312 (0.009) 0.366 (0.017)
Watermarking in Early vs Late Scales. We provide both qualitative and quantitative analysis for the scales in the Infinity architecture. We observe that the model produces high-level information in the image during the initial scales, forming a coarse shape of the main visual components, such as the object and the background (see Figure 4. In larger scales, the model predicts the details of the image (see Figure 5). As shown in Table 11, earlier and late scales contribute differently to the robustness of BitMark. Earlier scales are more robust against like CtrlRegen, as CtrlRegen does not change the content of the image itself.
Figure 4: Visualization of different images on different scales. Here, the image on scale $i$ refers to the cumulative image obtained by adding up $1$ to $i$ scales. We visualize five sets of images: 1) Original images without watermarking, 2) Watermarked (last 3), where the watermarks are generated on scale 11 to 13, 3) Differences (last 3), which is the difference between the original and watermarked images (last 3), 4) Watermarked (all), where the watermarks are generated on all scales, 5) Differences (all), which is the difference between the original and watermarked images (all). Here, we visualize scales 1-7.
Figure 5: Visualization of different images on different scales. Here, the image on scale $i$ refers to the cumulative image obtained by adding up $1$ to $i$ scales. We visualize five sets of images: 1) Original images without watermarking, 2) Watermarked (last 3), where the watermarks are generated on scale 11 to 13, 3) Differences (last 3), which is the difference between the original and watermarked images (last 3), 4) Watermarked (all), where the watermarks are generated on all scales, 5) Differences (all), which is the difference between the original and watermarked images (all). Here, we visualize scales 8-13.
Table 11: Robustness of BitMark applied to different scales with $\delta=2$. We report the TPR@1%FPR (%) for the different attacks on 5,000 watermarked images.
Scales Conventional Attacks Reconstruction Attacks
None Noise Blur Color Crop Rotation JPEG SD2.1-VAE CtrlRegen+
1-10 100.0 97.9 98.7 97.6 61.6 15.0 100.0 100.0 82.1
11-13 100.0 96.7 98.6 99.5 20.0 9.8 100.0 100.0 15.5
1-13 100.0 99.6 99.9 99.8 98.8 20.1 100.0 100.0 91.6
E.2 Image Quality
We analyze the impact of applying BitMark with different $\delta$ on the image quality. Figure 6 shows that applying BitMark with $\delta\leq 3$ keeps the semantic structure and quality of the image the same, while changing small features, e.g., the color of the train in the first row. For larger watermarking strengths, the image quality starts to deteriorate, where artifacts become visible in the final image.
Figure 6: Artifacts start to form in high detailed images already with $\delta=3$. Qualitative analysis of the impact of the choice of watermarking strength ($\delta$) on the perceived image quality.
E.3 Timing
We present the average elapsed time (in seconds) for the generation of an image with Infinity, when the watermark is applied to all scales (resolutions) and when using the standard generation process (without a watermark). For hardware information see Section C.2. Table 12 shows that applying BitMark results in negligible overhead to image generation and allows for a fast detection in under 0.5 seconds.
Table 12: Timing results for image processing and detection per image. We report the mean (std) for 1,000 images.
Setup Seconds/Image
BitMark 2.5817 (0.0063)
No watermark 2.3196 (0.0043)
Detection 0.4491 (0.0046)
E.4 Radioactivity
Table 13: We analyze the radioactivity of BitMark for class-conditional autoregressive models. We report the TPR@1%FPR (%).
Type of $M_{1}$ Type of $M_{2}$ Output of $M_{1}$ Output of $M_{2}$
Infinity-2B VAR-16 100.0 24.2
Infinity-2B VAR-20 100.0 25.8
Infinity-2B VAR-24 100.0 25.7
Infinity-2B VAR-30 100.0 25.6
Infinity-2B RAR-B 100.0 4.3
Infinity-2B RAR-L 100.0 3.3
Infinity-2B RAR-XL 100.0 3.9
Infinity-2B RAR-XXL 100.0 4.1
Table 13 reports the radioactivity of BitMark for the class-conditional autoregressive models, namely, VAR [32], which operates in the multi-scale manner similarly to Infinity, and RAR [37], which generates tokens only on a single scale in an autoregressive fashion. We find that BitMark exhibits detectable traces of radioactivity for the class-to-image autoregressive models models with TPR>>much-greater-than$>$ >1%@FPR=1%, with slightly lower traces for RAR (probably due to its single scale generation) than VAR. A potential cause of the overall relatively lower radioactivity for VAR and RAR is stemming from the class-conditional setting, instead of the text-to-image as in Infinity.
E.5 Rotation
Figure 7: BitMark is robust against rotation within ±$\pm 20$ degrees. The TPR@1%FPR for different $\delta$ and rotation degrees.
We ablate the robustness of BitMark against different degrees of rotation and different watermarking strengths $\delta$. We find that with $\delta=2$ BitMark is robust for rotation degrees ±$\pm 30°$. Additionally we find an increase in TPR@1%FPR for a rotation of 180°.
Appendix F Analysis for the Watermarks in the Sand Attack
As shown by the qualitative analysis in Figure 8, the watermarks in the sand attack causes a significant loss of details and produces many perceptible artifacts, limiting the applicability of this attack.
Figure 8: Watermarks in the Sand has non-negligible impact on image quality. Qualitative analysis of the impact of the watermarks in the sand. Here, we zoom in on some regions in the image to demonstrate the quality loss in details.
Appendix G Analysis for the Bit-Flipper Attack
Figure 9 shows different flipping strengths $\phi$ for $\delta=2$ on 1,000 watermarked samples. The most effective $\phi$ is $\phi=2.2$ with a TPR@1%FPR below 50% (see Table 14). The decrease in image quality for Bit-Flipper for different $\phi$ is shown in Figure 9, as well a more detailed analysis of $\phi=2.2$ on BitMark with a $\delta=2$ is displayed in Figure 10.
Figure 9: A higher flip factor leads to worse image quality but better watermark removal. Qualitative analysis of the impact of the Bit-Flipper with different $\phi$.
Figure 10: The Bit-Flipper has strong impact on the image quality. Here, we zoom in on some regions in the image to demonstrate the quality loss in details. $\phi=2.2$, $\delta=2$.
Table 14: Applying Bit-Flipper with different factors $\phi$ on watermarked images with $\delta=2$. We report the TPR@1%FPR for 1,000 Images.
$\phi$ TPR@1%FPR (%) $z$
1.8 0.878 10.06 (4.24)
2.0 0.845 9.17 (4.16)
2.2 0.472 5.52 (5.07)
2.4 0.595 8.06 (7.28)
2.6 0.52 6.97 (6.96)
Appendix H Limitations and Broader Impact
Limitations. While BitMark are robust against moderate rotations (99.9% TPR@1%FPR within rotation angles -20° to 20°), the detection accuracy degrades with larger rotation angles, dropping to 56.8% TPR@1%FPR for ±50° rotations. Although open-source methods have a strong capability for rotation correction, this can be a limitation for real-world scenarios. This rotation sensitivity, however, is an inherent issue in Infinity’s data augmentation pipeline, which makes the autoencoder less robust to large rotations. Moreover, the soft biasing parameter $\delta$ requires tuning, as values exceeding $\delta=3$ lead to noticeable quality degradation (FID increases from 29.61 to 127.44 when $\delta$ increases from 3 to 5), constraining the maximum achievable watermark strength. Although we evaluate the detection BitMark with a commonly adopted metric, i.e., true positive rate at very low false positive rates (1%), it could potentially become problematic in extreme-scale applications involving trillions of samples, where even small false positive rates might impact training data quality.
Broader Impact. Image generative models have the capabilities of generating photorealistic images, which are near imperceptible to human generated content. BitMark provides a method to watermark the content, without harming image generation capabilities of the model and allows for provenance tracking of generated images for models hosted under a private API. It supports the defense against model collapse, when training a model on data generated by itself.
Appendix I Additional Related Work
LLM Watermarking. Here, we review more related work regarding the LLM watermarking, which we build upon in our method. Beyond the soft biasing approach proposed by KGW [14], SynthID-Text [5] introduces tournament sampling during the model inference to enhance watermark detectability. This sampling strategy modifies the token sampling process to embed watermark signals. However, tournament sampling is only applicable to autoregressive models with large vocabulary sizes, making it incompatible with the binary bit sampling process used in Infinity. There are also works that incorporate watermarks by training or finetuning the models. For example, REMARK-LLM trains a learning-based encoder for watermark insertion and a decoder for watermark extraction, while using a reparameterization module to bridge the gap between the token distribution and the watermark message [39]. Moreover, potential malicious transformations are incorporated into the training phase to enhance the robustness of the watermarking. Although effective, these training-based methods require substantial computational resources and time. Therefore, we focus on inference-time watermarking methods that can be applied to Infinity without any retraining.
Sander et al. [28] investigate the radioactivity of LLM watermarking, which is the ability of watermarks to transfer across different model architectures. They design a specific protocol for amplifying the watermark signals after transferring across models. The radioactivity of watermarks also falls into the scope of our work, as our proposed watermark exhibits significant radioactivity across various model architectures. Regarding the undetectability of the watermark, recent studies show that the biases between watermarked and non-watermarked outputs from the LLM are identifiable with carefully designed prompts [17]. As a specialized application of LLM watermarking, Li et al. [15] design a watermarking scheme tailored for quantized LLMs. The proposed watermarks are only detectable on quantized models while remaining invisible for full-precision models.
Model Collapse. With the fraction of artificial generated content within the World Wide Web increasing, Shumailov et al. [31] investigate the phenomenon of model collapsing, which is defined as the decrease of generation quality when models are recursively trained on (partly) generated data. This decrease in generation quality is mainly characterized in overestimating probable events and underestimating low-probability events, as well as recursively adding noise (i.e., errors) which is generated during earlier model generations. Hence, this generated data is polluting future datasets, decreasing the potential quality of successor models and therefore highlighting the necessity to distinguish human- from artificial-generated data. BitMark mitigates model collapsing by robustly watermarking generated images to exclude said marked images from future datasets.